OpenStudy (anonymous):

Write the quadratic function f(x) = 3x^2 - 6x + 9 in vertex form.

OpenStudy (mathmate):

3x^2-6x+9=0 3(x^2-2x) +9=0 3(x-1)^2+9-3=0 3(x-1)^2+6=0